# Early warning indicators via latent stochastic dynamical systems

### 30 Dec 2023 v2

@@ Abstract
Detecting early warning indicators for abrupt dynamical transitions in complex systems or high-dimensional observation data is essential in many real-world applications, such as brain diseases, natural disasters, financial crises, and engineering reliability. To this end, we develop a novel approach: the directed anisotropic diffusion map that captures the latent evolutionary dynamics in the low-dimensional manifold. Then three effective warning signals (Onsager-Machlup Indicator, Sample Entropy Indicator, and Transition Probability Indicator) are derived through the latent coordinates and the latent stochastic dynamical systems. To validate our framework, we apply this methodology to authentic electroencephalogram (EEG) data. We find that our early warning indicators are capable of detecting the tipping point during state transition. This framework not only bridges the latent dynamics with real-world data but also shows the potential ability for automatic labeling on complex high-dimensional time series.
@@ 1 Introduction
High-dimensional evolutionary processes exist in many real-world systems, such as time-varying patterns of protein folding and unfolding, climate change, and artificial intelligence-assisted disease diagnosis [1, 2, 3]. Among these, one fascinating problem is studying complex brain activities over time, which also poses a challenge regarding the limited high-quality labeled data and multi-scale neuron models at different levels. Problems arise from modeling brain dynamics in high-dimensional space due to the complexity of the human brain system. Most of the multi-scale data-driven modeling approaches build surrogate models for brain activities. Electroencephalogram (EEG) signals are collected by dozens of electrodes placed on the scalp. Considering the low-dimensional representation from observations in electrophysiological experiments, neuroscientists have been investigating the latent dynamics of the brain via various dimension reduction techniques. Moreover, it shows that dimension reduction techniques lead to significantly enhanced performance by removing multi-collinearity and avoiding the curse of dimensionality.
@@
One methodology for learning low-dimensional latent dynamics comes from the variational auto-encoder framework, which ability to simultaneously reduce dimensionality while learning latent dynamics[4]. However, the disadvantages include the high computational cost, inaccuracy of reconstructed data, as well as weak explanation of effective coordinates of reduced systems. Therefore, other manifold learning methods are highly investigated to represent the dataset on a low-dimensional embedding space. Multiple studies consistently demonstrate that nonlinear manifold learning algorithms outperform linear ones. Among several effective nonlinear manifold learning methods, diffusion maps are widely used to analyze dynamical systems [5] and stochastic dynamical systems [6] on low-dimensional latent space, determine order parameters for chain dynamics [7], classify person from electrocardiogram recordings [8], and so on. Also, dynamics of brain activity from task-dependent functional magnetic resonance imaging data are modeled by a three-step framework including diffusion maps in [9]. Besides, the double diffusion maps on the latent space allow the approximation of the reduced dynamical models [10]. Further, the directed anisotropic diffusion map improves upon the isotropic diffusion by incorporating the local information. It considers the original data structure and adapts the diffusion process accordingly. Inspired by this and to efficiently distinguish between epilepsy-related data and pre-ictal data in the latent space, we add a drift function in the directed diffusion to construct an anisotropic diffusion map.
@@
After dimension reduction through manifold learning, the latent dynamics could be learned in either deterministic or stochastic ways. For example, Talmon et al. study the inference of latent intrinsic variables of dynamical systems from observed output signals [11]. Also, for partially observed data, Ouala et al. introduce a framework based on the data-driven identification of an augmented state-space model using a neural-network-based representation [12]. On the other hand, since stochastic differential equations are usually used to model nonlinear systems that involve uncertainty or randomness [13], it is of great interest to learn latent stochastic differential equations from real high dimensional data. To gain deeper insights into the latent characteristics, researchers frequently investigate the stochastic dynamical systems, which depend on the geometry and distribution of the latent space. For instance, Evangelou et al. combine diffusion map and Kramers–Moyal to construct a reduced, data-driven, parameter-dependent effective stochastic differential equation for electric-field mediated colloidal crystallization [14]. Besides, in general cases when the noise is non-Gaussian Lévy motion, there are various methods to realize the stochastic dynamical systems evolving in the latent space over time [15], as well as the corresponding transition dynamics [16, 17].
@@
In real world applications for complex disease diagnosis, early warning detection has attracted a great deal of attention for a long time. For instance, epilepsy is a neurological disease characterized by recurrent and unpredictable seizures, which are sudden bursts of abnormal brain activity. According to the World Health Organization, epilepsy affects around 50 million people worldwide, making it one of the most prevalent neural diseases. However, the mechanisms governing seizure activity are not well understood and many patients suffer a lot. Early detection and intervention of epileptic seizures can help patients receive timely and appropriate treatment, which might prevent or minimize the negative impacts of seizures on their health and quality of life. Liu et al. present a model-free computational method to detect critical transitions with strong noise by dynamical network markers [18]. Bury et al. develop a deep learning algorithm that provides early warning signals in systems by exploiting information of dynamics near tipping points [19]. In reality, the diagnosis of epilepsy requires experienced medical professionals. Diagnosis of epilepsy requires an experienced medical professional. Analyzing extensive EEG data to determine the onset of an epileptic seizure is a time-consuming and labor-intensive process. Additionally, relying on visual recognition can result in doctors inaccurately assessing the timing of the epilepsy occurrence. Therefore, in this paper, we propose to automatically detect the early warning signals from the latent stochastic dynamical system by introducing the Onsager-Machlup indicator, sample entropy, and transition probability.
@@
In summary, the goal of the present study is to model and analyze the brain activity data from epileptic patients to identify early warnings of epileptic seizures, with stochastic dynamical systems tools. Our main contributions are
@@
The remainder of this paper is structured as follows. In Section 2, we first design a drift function to obtain better low dimensional representation by the directed anisotropic diffusion map. Following that, we extract the latent stochastic dynamical system from high dimensional dataset, and establish specific early warning indicators. In Section 3, we validate our framework by numerical experiments on real EEG data. Finally, we summarize our results together with some conclusions in Section 4.
@@ 2.1 Anisotropic Diffusion Map
Let $X$ be the dataset in $\mathbf{R}^{n}$. Define a kernel $k$ on $X$ satisfying:
  • $k$ is symmetric, i.e. $k(x,y)=k(y,x)$,
  • $k$ is positivity preserving, i.e. $k(x,y)\geq 0$.
This kernel captures how closely the points in $X$ related to each other. It characterizes the connection between pairs of points within the symmetric graph, where the weight is determined by the kernel function $k$. The transition probabilities defined by the weight function in the graph directly reflect the local geometry. The technique to obtain the latent diffusion space is known as the normalized graph Laplacian construction[20]. Let $\mathcal{M}$ be a compact smooth manifold and $\mathcal{M}=\Phi(\mathcal{M})$, $\Phi:\mathcal{M}\rightarrow\mathbf{R}^{d}$ denote an isometric embedding into $d$-dimensional Euclidean space. Assume that the dataset $X$ is the entire manifold and $q(x)$ represents the density of the points on $\mathcal{M}$.
@@
The function $\nabla f$ defined on dataset can be obtained through derivation. To design a diffusion process adapted to the dataset, consider the directed kernel
$k_{\varepsilon}(x,y)=\exp(-\frac{\parallel x-y\parallel^{2}}{\varepsilon}-\frac{\parallel\nabla f(x)\cdot x-\nabla f(y)\cdot y\parallel^{2}}{\varepsilon^{2}}),$ (1)
@@
We adopt directed diffusion to show that the construction of diffusion kernel can be data-driven. The weight matrix $K$ corresponds to an anisotropic kernel on the observable space $X$, where the anisotropy is measured by the gradient $f$. By integrating over the second variable $k_{\varepsilon}$, the density $d_{\varepsilon}$ is represented by
@@
Dealing with finite data points, we should approximate integrals by finite sums and can describe the latent reduced order space by the spectral analysis. For most applications, we consider the bounded part will consist of a finite number of points. Here, the eigenvalues of the directed diffusion operator are distributed between 00 and $1$. In other words, we have $1=\lambda_{0}\geq\lambda_{1}\geq\lambda_{2}\geq\cdots$ with $\lambda_{i}\geq 0$. Since the first eigenvalue $\lambda_{0}=1$ and the eigenvector is a constant vector, the first $d$ eigenvalues starting from $\lambda_{1}$ and the corresponding eigenvectors capture the main features. The directed anisotropic diffusion map is given by $\Phi(x)=[\lambda_{1}\phi_{1}(x)\dots\lambda_{d}\phi_{d}(x)]^{T}$, where $\{\phi_{i}\}_{i=1}^{d}$ are eigenvectors. Several benefits are obtained by dimension reduction. It is well known that high dimensionality often degrades the dynamical performance, necessitating the use of lower-dimensional representations. Additionally, in the latent space, there is a significant reduction in the time required to train a model.
2.2 Extracting the Latent Stochastic Dynamical System
In order to depict the broader spectrum of nonlinear dynamics in the latent space, we proceed to the construction of the reduced order models. Here, the stochastic differential equations are preferred to carry the latent dynamics of the high-dimensional data. The drift term (vector field) characterizes the deterministic aspect of the system. While the diffusion term (fluctuation) pertains to the stochastic component of the system and depends on a random process, such as Brownian motion or Lévy process.
Let $z(t)$ be a stochastic variable governed by the stochastic differential equation
@@ 2.2 Extracting the Latent Stochastic Dynamical System
where $\delta_{B^{j}_{t}}=B_{t_{i+1}}-B_{t_{i}}\sim\mathcal{N}(0,\delta_{t})$, $j=1,\ldots,d$. Conditioned on $(z(t_{i}),\delta_{t})$, the point $z(t_{i+1})$ obeys a multivariate normal distribution
$z(t_{i+1})\sim\mathcal{N}(z(t_{i})+\delta_{t}\mu(z(t_{i})),\delta_{t}\eta(z(t_{i}))^{2}).$
Hence, the neural networks are trained by minimizing the loss function $\mathcal{L}(\theta)$, which is defined as
$\mathcal{L}(\theta)=\frac{(z(t_{i+1})-z(t_{i})-\delta_{t}\mu_{\theta}(z(t_{i}))^{2})}{\delta_{t}\eta_{\theta}(z(t_{i}))^{2}}+\log\mid\delta_{t}\eta_{\theta}(z(t_{i}))^{2}\mid+\log(2\pi).$ (5)
@@ 2.3 Early Warning Indicators
When existing latent variables in complex disease processes, a critical concern revolves around the concept of early warning indicators. To identify abrupt changes and issue early warnings, we introduce the indicators by Onsager-Machlup action functional, sample entropy, and transition probability.
@@ 2.3.1 Onsager-Machlup Indicator
Consider the following stochastic differential equation in the state space $\mathbb{R}^{m}$
@@
where $\mu:\mathbb{R}^{m}\to\mathbb{R}^{m}$, $B_{t}$ is a standard $m$-dimensional Brownian motion, and $\eta$ is a diagonal matrix with positive constants. Onsager and Machlup are the first to consider the probability of paths of a diffusion process as the starting point of a theory of fluctuations [23]. The outcome of a functional integral over the process paths is subsequently termed the Onsager-Machlup function, denoted by
@@
where $\psi$ is a curve in $\mathbb{R}^{m}$.
Let time series $z(t)$ be a stochastic variable governed by the stochastic differential equation in the latent space. Recalling Sec. 2.2, two neural networks $\mu_{\theta}$ and $\eta_{\theta}$ are constructed to estimate $\mu$ and $\eta$, whose inputs and weights are represented by the basis functions and coefficients, respectively. Here, we define the Onsager-Machlup indicator by
@@
where $\dot{z}$ is the derivative function written as
$\dot{z}(s)=\frac{z(s+ds)-z(s-ds)}{2ds}.$
If there are notable variations in Onsager-Machlup indicator, we will issue an early warning.
@@ 2.3.2 Sample Entropy Indicator
Sample entropy, introduced by Richman and Moorman [24], serves as a statistical metric for quantifying the complexity of time series data. It is particularly effective in capturing self-similarity and irregularity, making it a robust tool for analyzing complex systems. For $d$ interdependent time series, we introduce the multidimensional sample entropy to measure the latent system complexity [25]. Specifically, it approximately equals to the negative natural logarithm of the conditional probability that two sub-sequences similar for consecutive data points remain similar for the next $p$ time points, where the sub-sequences can originate from either the same or different time series.
@@
If the superscripts $\alpha=\beta$, the subscripts $i\neq j.$ Assume that two vectors are close if their distance $d(Z^{\alpha}_{i}(m,q),Z^{\beta}_{j}(m,q))<r\max\{\sigma^{\alpha},\sigma^{\beta}\}$, where $r$ is the tolerance for accepting matches and $\sigma^{\alpha}$ and $\sigma^{\beta}$ are the standard deviations, allowing measurements on dataset with different amplitudes to be compared. Now we consider a set $A(m+p,q)$ of all vectors satisfying $Z^{\alpha}_{k}(m+p,q)\subset z^{\alpha}[t_{1}:t_{l}]$ and $\alpha=1,\ldots,d$. Note that the maximum of $k$ in $A(m+p,q)$ is $n_{z}$. To make the number of vectors of length $m$ equal to that of length $m+p$, the set $B(m,q)$ is characterized as $\{Z^{\alpha}_{k}(m,q):Z^{\alpha}_{k}(m,q)\subset z^{\alpha}[t_{1}:t_{l}],1\leq\alpha\leq d,1\leq k\leq n_{z}\}$. Define $B(m,q,r,l)$ as the number of close vectors in $B(m,q)$. Similarly, denote the number of close vectors in $A(m+p,q)$ by $A(m+p,q,r,l)$. Hence, the sample entropy for multidimensional system is represented by
$SampEn(m,p,q,r,l)=-\log\left(\frac{A(m+p,q,r,l)}{B(m,q,r,l)}\right),$
in the time interval $[t_{1},t_{l}]$. The parameters $l,\>m,\>p,\>q$ and $r$ must be fixed for each calculation. There is no confusion about the parameters, so in time window $[t-l+1,t]$
@@
In the event of significant fluctuations in sample entropy, we will give an early warning.

### 05 Apr 2024 v3

@@ Abstract
Detecting early warning indicators for abrupt dynamical transitions in complex systems or high-dimensional observation data is essential in many real-world applications, such as brain diseases, natural disasters, and engineering reliability. To this end, we develop a novel approach: the directed anisotropic diffusion map that captures the latent evolutionary dynamics in the low-dimensional manifold. Then three effective warning signals (Onsager-Machlup Indicator, Sample Entropy Indicator, and Transition Probability Indicator) are derived through the latent coordinates and the latent stochastic dynamical systems. To validate our framework, we apply this methodology to authentic electroencephalogram (EEG) data. We find that our early warning indicators are capable of detecting the tipping point during state transition. This framework not only bridges the latent dynamics with real-world data but also shows the potential ability for automatic labeling on complex high-dimensional time series.
@@ 1 Introduction
High-dimensional evolutionary processes exist in many real-world systems, such as time-varying patterns of protein folding and unfolding, climate change, and artificial intelligence-assisted disease diagnosis [1, 2, 3]. Among these, one fascinating problem is studying complex brain activities over time. Due to the complexity of the human brain system, most of the multi-scale data-driven modeling approaches build surrogate models for brain activities. Electroencephalogram (EEG) signals are collected by dozens of electrodes placed on the scalp. Considering the low-dimensional representation from observations in electrophysiological experiments, neuroscientists have been investigating the latent dynamics of the brain via various dimension reduction techniques, leading to significantly enhanced performance by removing multi-collinearity and avoiding the curse of dimensionality.
@@
One methodology for learning low-dimensional latent dynamics comes from the variational auto-encoder framework, which is able to simultaneously reduce dimension while learning latent dynamics[4]. However, the disadvantage is the weak explanation of effective coordinates of reduced systems. Therefore, other dimension reduction methods are highly investigated to represent the geometric structure of the dataset on a low-dimensional space [5]. Among the nonlinear methods, diffusion maps [6] are widely used to analyze dynamical systems [7] and stochastic dynamical systems [8] on low-dimensional latent space, determine order parameters for chain dynamics [9], classify person from electrocardiogram recordings [10], and so on. Also, dynamics of brain activity from task-dependent functional magnetic resonance imaging data are modeled by a three-step framework including diffusion maps in [11]. Furthermore, the directed anisotropic diffusion map [12] improves upon the isotropic diffusion by incorporating the local information. It considers the original data structure and adapts the diffusion process accordingly via a directed term. To efficiently distinguish between epilepsy-related data and pre-ictal data in the latent space, we adopt the data-driven directed diffusion to construct an anisotropic diffusion map.
@@
After dimension reduction through diffusion map, the latent dynamics could be learned in either deterministic or stochastic ways. For example, Talmon et al. study the inference of latent intrinsic variables of dynamical systems from observed output signals [13]. Also, for partially observed data, Ouala et al. introduce a framework based on the data-driven identification of an augmented state-space model by a neural network [14]. On the other hand, since stochastic differential equations are usually used to model nonlinear systems that involve uncertainty or randomness [15], it is of great interest to learn latent stochastic differential equations from real high dimensional data. To gain deeper insights into the latent characteristics, researchers frequently investigate the stochastic dynamical systems, which depend on the geometry and distribution of the latent space. For instance, Evangelou et al. combine diffusion map and Kramers–Moyal to construct an effective stochastic differential equation for electric-field mediated colloidal crystallization [16]. Besides, in general cases when the noise is non-Gaussian, there are various methods to realize the stochastic dynamical systems evolving in the latent space over time [17], as well as the corresponding transition dynamics [18, 19].
@@
In real-world applications for complex disease diagnosis, early warning detection has attracted a great deal of attention. For instance, epilepsy is a neurological disease characterized by recurrent and unpredictable seizures, which are sudden bursts of abnormal brain activity. According to the World Health Organization, epilepsy affects around 50 million people worldwide. Early detection and intervention of epileptic seizures can help patients receive timely and appropriate treatment, which might prevent or minimize the negative impacts of seizures on their health and quality of life. Liu et al. present a model-free computational method to detect critical transitions with strong noise by dynamical network markers [20]. Bury et al. develop a deep learning algorithm that provides early warning signals in systems by exploiting information on dynamics near tipping points [21]. In reality, the diagnosis of epilepsy requires experienced medical professionals. Analyzing extensive EEG data to determine the onset of an epileptic seizure is a time-consuming and labor-intensive process. Additionally, relying on visual recognition can result in doctors inaccurately assessing the timing of the epilepsy occurrence. Therefore, in this paper, we propose to automatically detect the early warning signals from the latent stochastic dynamical system by introducing the Onsager-Machlup indicator, sample entropy indicator, and transition probability indicator.
@@
In summary, the goal of the present study is to model and analyze the brain activity data from epileptic patients to identify early warnings of epileptic seizures automatically, with stochastic dynamical systems tools. Our main contributions are
@@
The remainder of this paper is structured as follows. In Section 2, to obtain a better low dimensional representation, we adopt the directed anisotropic diffusion map with a data-driven directed term. Following that, we extract the latent stochastic dynamical systems from the high dimensional dataset, and establish specific early warning indicators. In Section 3, we validate our framework by numerical experiments on real EEG data. Finally, we summarize in Section 4.
@@ 2.1 Anisotropic Diffusion Map
Let $X$ be the dataset in $\mathbf{R}^{n}$, and define a kernel $k(x,y)$ on $X\times X$. Note that $k(x,y)=1$ if and only if $x=y$, and $k(x,y)$ has exponential decay with distance. This kernel captures some geometric features from the data and the probability of transitions between $x$ and $y$. The technique to obtain the latent diffusion space is known as the normalized graph Laplacian construction[22]. Assume that a set of points on or near a $d$-dimensional manifold are embedded in a high-dimensional space $\mathbf{R}^{n}$, and the manifold exists. The map $\Phi:\mathcal{M}\rightarrow\mathbf{R}^{d}$ learns the geometric structure of the data $X$ sampled from $\mathcal{M}$ with the density $q$.
@@
To design a diffusion process adapted to the dataset, consider the directed kernel, mentioned in [12, 23],
${k_{\varepsilon}(x,y)=\exp(-\frac{\parallel x-y\parallel^{2}}{\varepsilon}-\frac{\parallel\langle\nabla_{x}f,x-y\rangle\parallel^{2}}{\varepsilon^{2}}),}$ (1)
@@
In the later experiment, we will show the benefit of anisotropic directed diffusion map, since for real data applications, it is more capable of capturing the intrinsic distribution shift of the data, as a return, generating more effective early warning indicators from the low-dimensional latent dynamics.
The weight matrix $K$ corresponds to an anisotropic kernel on the observable space $X$. By integrating over the second variable $k_{\varepsilon}$, the density $d_{\varepsilon}$ is represented by
@@
Dealing with finite data points, we should approximate integrals by finite sums and can describe the latent reduced order space by the spectral analysis. For most applications, we consider the bounded part will consist of a finite number of points. Here, the eigenvalues of the directed diffusion operator are distributed between 00 and $1$. In other words, we have $1\geq\lambda_{1}\geq\lambda_{2}\geq\cdots$. The first $d$ eigenvalues and the corresponding eigenvectors capture the main features. The directed anisotropic diffusion map is given by $\Phi(x)=[\lambda_{1}\phi_{1}(x)\dots\lambda_{d}\phi_{d}(x)]^{T}$, where $\{\phi_{i}\}_{i=1}^{d}$ are eigenvectors. Several benefits are obtained by dimension reduction. It is well known that high dimensionality often degrades the dynamical performance, necessitating the use of lower-dimensional representations. Additionally, in the latent space, there is a significant reduction in the time required to train a model.
2.2 Extracting Latent Stochastic Dynamical System
In order to depict the broader spectrum of nonlinear dynamics in the latent space, stochastic differential equations are preferred to carry the low dimensional latent dynamics of the high-dimensional data. The drift term (vector field) characterizes the deterministic aspect of the system. While the diffusion term (fluctuation) pertains to the stochastic component of the system under noise such as Brownian motion or Lévy process.
Let $z(t)$ be a stochastic process governed by the stochastic differential equation
@@ 2.2 Extracting Latent Stochastic Dynamical System
where $\delta_{B^{j}_{t}}=B_{t_{i+1}}-B_{t_{i}}\sim\mathcal{N}(0,\delta_{t})$, $j=1,\ldots,d$. Conditioned on $(z(t_{i}),\delta_{t})$, the point $z(t_{i+1})$ obeys a multivariate normal distribution $\mathcal{N}(z(t_{i})+\delta_{t}\mu(z(t_{i})),\delta_{t}\eta^{2}(z(t_{i}))).$ Hence, the neural networks are trained by minimizing the loss function $\mathcal{L}(\theta)$, which is defined as
$\mathcal{L}(\theta)=\frac{(z(t_{i+1})-z(t_{i})-\delta_{t}\mu_{\theta}(z(t_{i})))^{2}}{\delta_{t}\eta^{2}_{\theta}(z(t_{i}))}+\log\mid\delta_{t}\eta^{2}_{\theta}(z(t_{i}))\mid+\log(2\pi).$ (5)
@@ 2.3 Early Warning Indicators
After obtaining the latent stochastic dynamical system from high dimensional data, we can now identify some abrupt critical transitions, through indicators created by Onsager-Machlup action functional, sample entropy, and transition probability.
@@ 2.3.1 Onsager-Machlup Indicator
Consider the following stochastic differential equation in the state space $\mathbb{R}^{d}$
@@
where $\mu:\mathbb{R}^{d}\to\mathbb{R}^{d}$, $B_{t}$ is a standard $d$-dimensional Brownian motion, and $\eta$ is a diagonal matrix with positive constants. Onsager and Machlup are the first to consider the probability of paths of a diffusion process as the starting point of a theory of fluctuations [26]. The outcome of a functional integral over the process paths is subsequently termed the Onsager-Machlup function, denoted by
@@
where $\psi$ is a curve in $\mathbb{R}^{d}$.
Let time series $z(t)$ be a stochastic variable governed by the stochastic differential equation in the latent space. Recalling Sec. 2.2, two neural networks $\mu_{\theta}$ and $\eta_{\theta}$ are constructed to estimate $\mu$ and $\eta$, whose inputs and weights are represented by the basis functions and coefficients, respectively. Here, we define the Onsager-Machlup functional by
@@
where $\dot{z}$ is the derivative function written as $\dot{z}(s)\approx\frac{z(s+ds)-z(s-ds)}{2ds}$ with the time step $ds$.
@@ 2.3.2 Sample Entropy Indicator
Sample entropy [27], serves as a statistical metric for quantifying the complexity of time series data. It is particularly effective in capturing self-similarity and irregularity, making it a robust tool for analyzing complex systems. For $d$ interdependent time series, we introduce the multidimensional sample entropy to measure the latent system complexity [28]. Specifically, it approximately equals to the negative natural logarithm of the conditional probability that two sub-sequences similar for consecutive data points remain similar for the next $p$ time points.
@@
If the superscripts $\alpha=\beta$, the subscripts $i\neq j.$ Assume that two vectors are close if their distance $d(Z^{\alpha}_{i}(m,q),Z^{\beta}_{j}(m,q))<r\max\{\sigma^{\alpha},\sigma^{\beta}\}$, where $r$ is the tolerance for accepting matches and $\sigma^{\alpha}$ and $\sigma^{\beta}$ are the standard deviations, allowing measurements on dataset with different amplitudes to be compared. Now we consider a set $A(m+p,q)$ of all vectors satisfying $Z^{\alpha}_{k}(m+p,q)\subset z^{\alpha}[t_{1}:t_{l}]$ and $\alpha=1,\ldots,d$. Note that the maximum of $k$ in $A(m+p,q)$ is $n_{z}$. To make the number of vectors of length $m$ equal to that of length $m+p$, the set $B(m,q)$ is characterized as $\{Z^{\alpha}_{k}(m,q):Z^{\alpha}_{k}(m,q)\subset z^{\alpha}[t_{1}:t_{l}],1\leq\alpha\leq d,1\leq k\leq n_{z}\}$. Define $B(m,q,r,l)$ as the number of close vectors in $B(m,q)$. Similarly, denote the number of close vectors in $A(m+p,q)$ by $A(m+p,q,r,l)$. Hence, the sample entropy $SampEn(m,p,q,r,l)$ for multidimensional system is represented by $-\log\left(\frac{A(m+p,q,r,l)}{B(m,q,r,l)}\right)$ in the time interval $[t_{1},t_{l}]$. The parameters $l,\>m,\>p,\>q$ and $r$ must be fixed for each calculation. So in time window $[t-l+1,t]$, we have the entropy indicator